How do I write an equation of the line that passes through (1,6) and is parallel to the line x-y=4 ?

Answer by stanbon(75887)

How do I write an equation of the line that passes through (1,6) and is parallel to the line x-y=4 ?
-----
Find the slope of the given line::
y = x-4
slopw = 1
-----
Lines parallel to the given line must have slope = 1.
-----
Form: y = mx + b
Solve for "b":
6 = 1*1 + b
b = 5
------
Equation:
y = x + 5
--------------------
Cheers,
Stan H.
=====================